OpenStudy (anonymous):

Write the next term of the sequence: 1/2, 4/3, 9/7, 49/16, ...

OpenStudy (anonymous):

try figuring out what is the difference between 49/16 and 9/7, then 9/7 with 4/3, then 4/3 and 1/2

terenzreignz (terenzreignz):

This is complicated, but I have reason to believe that the answer is 256/32 = 8.

terenzreignz (terenzreignz):

Look, after 1/2, the numerators are the squares of the PREVIOUS denominator. That's how I got 256. And the denominators after 1/2, are always the previous denominator plus a perfect square... like this : 3 = 2 + 1 7 = 3 + 4 16 = 7 + 9 So I figured the next denominator must be the previous denominator, 16, plus the next perfect square, which also happened to be 16, so.. 32 = 16 + 16

OpenStudy (anonymous):

saw that

OpenStudy (anonymous):

so the next demoninator should be 49+16

OpenStudy (anonymous):

but I don't see any holonomic sequence

OpenStudy (anonymous):

me too I do not see a sequence... I am having a hard time solving this.

OpenStudy (anonymous):

I see the denominator but not the numerator

OpenStudy (anonymous):

I get it now

OpenStudy (anonymous):

the numerator is the square of the denominator

OpenStudy (anonymous):

next is 256/65

OpenStudy (anonymous):

saw it! thanks!

terenzreignz (terenzreignz):

@panlac01 Why is the next denominator 65?

OpenStudy (anonymous):

the denominator is the sum of the numerator and denominator of the previous fraction except 1/2

OpenStudy (anonymous):

49+16

terenzreignz (terenzreignz):

Makes sense... Far simpler

OpenStudy (anonymous):

lol

terenzreignz (terenzreignz):

I tried my best to make sense of it.. it looks like I'm complicating things too much. Thanks!

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

np. no textbook can teach you this, so I answered to my fullest.

OpenStudy (anonymous):

haha!

terenzreignz (terenzreignz):

I wasn't expecting to learn this through textbooks... this is intuition :D

OpenStudy (anonymous):

:)